# Scaling Up On-Device LLMs via Active-Weight Swapping Between DRAM and Flash

### v1

@@ 1. Introduction
Adaptive DRAM usage has been previously investigated for traditional non-autoregressive DNNs (e.g., CNN and Bert) through DRAM-Flash swapping (Li et al., 2024; Yi et al., 2023). However, the fundamental difference in workload characteristics hinders the direct application of these methods to LLMs. Existing techniques rely on the computation-intensive feature of traditional DNNs, so the current operator computation can overlap the loading of the next operator. While this overlap is present in the LLM prefilling stage, the significantly more time-consuming autoregressive decoding phase is bottlenecked by memory access. Consequently, realizing user-oblivious adaptive memory management for LLM inference necessitates minimizing Flash data loading to mitigate the substantial disparity between memory and Flash bandwidth (5$\sim$ 5$\times$ 5 × on mobile phones).
@@
Figure 1. The perplexity of LLaMA-3-8B model on various memory cost. Ours shows the best performance compared with SOTA quantization (Q), pruning (P) and contextual sparsity (SP). Each point on the scaling line means a sparsity ratio.
@@
Several works have explored contextual sparsity(Liu et al., 2023; Xue et al., 2024; Song et al., 2024b; Wang et al., 2024b; L et al., 2024; Mirzadeh et al., 2023; Song et al., 2024c, a), but gaps remain in addressing the challenges above. Some methods like Deja Vu (Liu et al., 2023), PowerInfer (Song et al., 2024b) and LLM in a flash (Alizadeh et al., 2023) use available ReLU-based models to generate zero activations and introduce additional predictors (GB memory cost) to forecast these zeros. However, modern LLMs used in productions (e.g., Llama) rarely use ReLU-based architecture due to its inferior accuracy (Touvron et al., 2023) (see Fig.14(b)). There are also works performing continued pre-training to transform available models to ReLU or ReLU-variant based, such as PowerInfer-2 (Xue et al., 2024), TurboSparse (Song et al., 2024c), ProSparse (Song et al., 2024a), and Q-Sparse (Wang et al., 2024b). These works require training on hundreds of billions of tokens and consume substantial hardware resources. TEAL (Liu et al., 2025a) proposes a training-free, magnitude-based sparsity method (see Fig. 3), where only activations above a threshold are computed. However, the active weights cannot be predicted, but only be identified after the input activation is ready. Additionally, the method is empirical, and there is no mechanism to compensate for the accuracy loss due to the potential misidentification of active weights. Therefore, current techniques fall short of achieving adaptive memory usage for LLMs.
@@
Firstly, Cross-layer active weight preloading. To address the sequential dependency issue of active weights with its input activation in order to enable computation and loading overlapping, we propose cross-layer active weight preloading. It creatively utilizes the current layer’s activation to pre-identify the next n layers’ active weights. It is based on the key obeservation that due to the widely used residual connection, the activation magnitude distribution across layers share significant similarity (¿80% shown in Fig. 4(a)). For the active weights that missed by pre-loading, ActiveFlow loads on-demand when the actual activation is ready.
@@
We implement ActiveFlow and evaluate it on different mobile phones (OnePlus 12, Pixel 6, and Infinix Zero). Results (Fig. 1, more in Sec. 7) show that ActiveFlow achieves the inference performance-cost Pareto frontier among existing efficiency optimization methods, including state-of-the-art quantization (DB-LLM (Chen et al., 2024) and PB-LLM (Shang et al., 2023)), pruning (CPSP (Wang et al., 2024d) and RIA (Yi and Richtárik, 2025)), and contextual sparsity (Teal (Liu et al., 2025b)), demonstrating its practical value. Particularly, under the same model quality and speed, ActiveFlow reduces the DRAM usage by up to 40% for Llama 7B compared to llama.cpp. Under the same sparsity ratio, ActiveFlow can reduce memory by 2$\times$ compared to Teal. ActiveFlow is the first to successfully deploy the original Mixtural-8x7B 4bit model (Jiang et al., 2024) (no ReLU introduced) on a mid-range pixel-6 phone, achieveing 1.8 tokens/s with 2.9 GB memory cost.
@@ 4. Active Weight Swapping Pipeline
$sp$ sparsity of LLM
$hr$ average hit rate of weight cache
$si$ average similarity of cross-layer group
@@
$N$ Layer number of a cross-layer group
@@ 5. Self-Distillation for Top-K Sparse LLM
Even with superior quality compared to other sparsity techniques, Top-K activation sparsity still introduce approximation for active weights selection. To compensate this, inspired by quantization-aware self-distillation, we propose Top-K sparsity-aware self-distillation. It has been integrated with the existing quantization self-distillation framework, together with the customer data (e.g., agent) fine-tuning as a standard step for actual LLM deployment on device, without adding extra overhead.
The reason to choose self-distillation is because traditional supervised fine tuning often fails to recover the model’s nuanced weight distributions and activation patterns, leading to performance loss. By comparison, self-distillation uses the full model’s output distribution as a soft target to preserve essential distributional details and weight correlations, thereby reducing errors and maintaining fine inference details for more robust, generalizable performance. It boosts accuracy with only a few to tens of GPU hours to learn on a few thousand samples, and generalizes across various sparsity levels.
Standard self-distillation faces two major challenges to apply for Top-K sparsity. First, gradient vanishing occurs because the sparsity mask introduces many zero values, causing most backpropagated gradients to also become zero, which slows down training. Second, loss mismatch arises because traditional cross-entropy loss does not effectively capture the relationship between the sparse LLM and the original model, leading to suboptimal distillation results.
Figure 13. The forward and backward process of self-distillation between Top-K activation sparsity and full weight model.
To overcome these issues, we propose top-k sparsity self-distillation as illustrated in Fig. 13, which introduces two key improvements:
The first is gradient straight-through estimation, a technique that reduces the impact of zero-masked values on gradient propagation, allowing more gradients to pass through and accelerating convergence. The second is the use of combine KLD(Kullback-Leibler Divergence)-CE(cross-entropy) loss, which better aligns the sparse LLM’s loss with that of the original model, improving the effectiveness of the distillation process.
5.1. Gradient STE and SD loss
Gradient Straight-Through Estimation. Sparsity leads to a large number of zero elements, which causes the gradients to be unable to update during backpropagation. To mitigate this, gradient straight-through estimation (STE) is employed. As shown in Equations 10 and 11, STE preserves the learning signal by replacing the gradient of the masking operation with an identity function during the backward pass:
@@ 5.1. Gradient STE and SD loss
By doing so, STE effectively bypasses the zero-masking effect in the gradient computation, allowing the gradients to flow as if the masking were transparent. This ensures that even in the presence of high sparsity, the model receives sufficient gradient information to update its parameters. Consequently, STE not only accelerates convergence by preventing gradient decay but also enhances the robustness of sparse inference training, thereby enabling the model to better retain and adapt the critical weight distributions and activation patterns.
Sparse-Distillation Loss. In our self-distillation framework, we denote $P_{T}$ as the output probability distribution of the teacher model and $P_{S}$ as that of the student model. We first define the Kullback-Leibler divergence as:
(12) $\mathcal{D}_{\text{KL}}(P\parallel Q)$ $=\sum_{i}P(i)\log\frac{P(i)}{Q(i)}$
In addition, we incorporate the cross-entropy loss between the teacher-provided labels $y_{T}$ and the student outputs $y_{S}$, denoted as $CE(y_{T},y_{S})$, which measures the discrepancy between the student prediction distribution and the teacher’s soft labels. Combining these two components, the self-distillation loss is formulated as:
(13) $\mathcal{L}_{\text{SD}}$ $=\gamma\,\mathcal{D}_{\text{KL}}(P_{T}\parallel P_{S})+(1-\gamma)\,CE(y_{T},y_{S}),$
where $\gamma\in[0,1]$ is a weighting factor that balances the two loss terms. It depends on sparsity ratio. Under high sparsity, $\gamma$ tends toward 0. Under low sparsity, it comes to 1. This occurs because, at high sparsity levels, sparse model’s output distribution deviates significantly from the original model. As a result, using KL divergence to measure the discrepancy becomes less reliable and may lead to performance degradation. In these cases, employing cross-entropy loss, which directly compares the predicted labels, provides a more effective training signal.
Minimizing $\mathcal{L}_{\text{SD}}$ encourages the student model not only to mimic the overall distribution of the teacher model but also to closely match its output values. This approach preserves crucial distributional information and weight correlations even under conditions of quantization and sparsity.
5.2. One-Distill-All-Scale: A Key Advantage of Distillation
A major advantage of our self-distillation approach is its capability of one-distill-all-scale. In conventional fine-tuning or distillation procedures, separate processes are often required for models at different sparsity levels, which can be both time-consuming and computationally expensive. In contrast, our method performs a single distillation at a high sparsity level, capturing the essential distributional characteristics and weight correlations of the original model.
Thanks to this process, the distilled student model can be effectively applied to scenarios with lower sparsity without additional fine-tuning. This is because the self-distillation not only transfers knowledge from the full teacher model but also establishes a robust framework that generalizes well across different sparsity scales. As a result, the model maintains consistent performance and efficiency, drastically reducing the overall training and deployment costs.
In summary, the one-distill-all-scale property enables our approach to streamline the distillation process and ensure adaptability across a range of sparsity levels, making it particularly attractive for efficient large-scale model inference.
@@ 6. Implementation
Self-distillation. In order to implement and test sparsity-aware self-distillation, we extend BitDistiller (Du et al., 2024) into an open-source framework called Sparse-Distillation. We have developed a plug-and-play module that enables sparse model inference and training with just one line of code. This framework also supports quantization-aware sparsification. Specifically, we add a predictor to each block and perform activation sparsification by loading pre-generated block-level thresholds. During backpropagation, we incorporate a gradient straight-through estimation (STE) layer, and in the loss function design, we provide an SDloss option. In our self-distillation experiments, we train on a mixed dataset from alpaca-wikitext-c4, with each epoch containing approximately 9k data samples. A full training run comprises five epochs, with a learning rate of $8\times 10^{-6}$ and 4-bit quantization.
@@ 7. Evaluation
7.1. Evaluation setup
Measurement. Our evaluation focuses on decoding speed, perplexity, accuracy, latency, hit rate, memory cost, power, and energy consumption. We use the clock_gettime function to record start and end timestamps, computing latency as the difference between them. We measure the total number of decoded tokens and the total decoding time, calculating speed as $N_{tokens}/Latency$. We use lm-eval-harness, a widely used LLM evaluation framework, to measure perplexity on the WikiText-2 tasks. We track cache hits and misses, computing the hit rate as $N_{hit}/(N_{hit}+N_{miss})$. We analyze memory cost using the Android Studio Profiler. We obtaine current and voltage values by reading system files (voltage_now and current_now) to calculate power consumption. These values are collected every 0.5 seconds on average, and we use the decoding latency to compute the overall energy consumption.
@@ 7.2. End-to-end performance
Perplexity. We evaluate the perplexity and accuracy of different LLMs under varying memory cost conditions as illustrated in Fig. 14(b). For the LLaMA-2-7B model and LLaMA-3-8B model, our method achieved results comparable to the full-weight memory setting at 60% memory cost. For Mixtral-8x7B, we achieve a perplexity comparable to the baseline (24.6GB) while using only 4.4GB of memory. With further memory reduction, perplexity increases due to the increased sparsity of the LLM, which leads to a decline in inference capability.
@@ 7.3. Technique breakdown
Cross-layer-group pipeline. First, we examine the effect of the cross-layer-group pipeline on decoding speed, as shown in Fig. 15. We used a 60% sparsity LLaMA-2-7B model and tested it across three devices. Our baseline consisted of serial computation and memory reads. Experimental results show that when the layer number in a cross-layer group is set to 1, the average speedup across all three devices is 10%. However, increasing the layer number to 4 results in a 120% performance improvement, as it enhances the efficiency of flash memory reads. Finally, with the addition of Dynamic Cache, our method achieves 2$\times$, 2.3$\times$, and 3$\times$ speedups over the baseline on the three devices, respectively.
@@
Contextual caching policy. In Fig. 17, we compare the performance of context-level cache and task-level cache, where the task-level cache is built based on hot weight statistics from the WikiText-2 dataset. Fig. 17(a) presents the cache hit rate of context-level and task-level caches under different token lengths, evaluated on a sequence of BoolQ dataset. When the token length is 10, the context-level cache achieves a 13% higher hit rate than the task-level cache. As the token length increases to 40, the context-level cache’s hit rate drops from 77% to 74%, yet it remains 10% higher than the task-level cache. Fig. 17(b) shows the cache hit rate across different downstream tasks. The task-level cache’s hit rate varies between 54% and 74% across different datasets due to variations in hot weight distributions among tasks. In contrast, our context-level cache effectively adapts to these variations, achieving an average hit rate improvement of 12% over the task-level cache.
Figure 17. The performance of task-level and context-level cache.
Figure 18. The perplexity of baseline and top-k sparsity-aware self-distillation.
@@
Top-k sparsity-aware self-distillation. Fig. 18 shows the perplexity changes of baseline and our top-k sparsity-aware self-distillation. Compared with the baseline, our approach substantially lowers the perplexity, especially when the sparsity exceeds 80%, resulting in remarkable improvements in LLM accuracy. The improvement in accuracy can be attributed to the incorporation of the gradient straight-through estimator and the adaptive combined loss function in our self-distillation process.

### v2

@@ 1. Introduction
Adaptive DRAM usage has been previously investigated for traditional non-autoregressive DNNs (e.g., CNN and Bert) through DRAM-Flash swapping (Li et al., 2024; Yi et al., 2023). However, the fundamental difference in workload characteristics hinders the direct application of these methods to LLMs. Existing techniques rely on the computation-intensive feature of traditional DNNs, so the current operator computation can overlap the loading of the next operator. While this overlap is present in the LLM prefilling stage, the significantly more time-consuming autoregressive decoding phase is bottlenecked by memory access. Consequently, realizing user-oblivious adaptive memory management for LLM inference necessitates minimizing Flash data loading to mitigate the substantial disparity between memory and Flash bandwidth ($\sim 5\times$ on mobile phones).
@@
Figure 1. The perplexity versus cost of LLaMA-3-8B model. Ours shows the Pareto frontier compared with SOTA model compression methods including quantization (Q), pruning (P) and contextual sparsity (SP). Each point on the scaling line means a sparsity ratio.
@@
Several works have explored contextual sparsity(Liu et al., 2023; Xue et al., 2024; Song et al., 2024b; Wang et al., 2024b; L et al., 2024; Mirzadeh et al., 2023; Song et al., 2024c, a), but gaps remain in addressing the challenges above. Some methods like Deja Vu (Liu et al., 2023), PowerInfer (Song et al., 2024b) and LLM in a flash (Alizadeh et al., 2023) use available ReLU-based models to generate zero activations and introduce additional predictors (GB memory cost) to forecast these zeros. However, modern LLMs used in productions (e.g., LLaMA) rarely use ReLU-based architecture due to its inferior accuracy (Touvron et al., 2023) (see Fig.14(b)). There are also works performing continued pre-training to transform available models to ReLU or ReLU-variant based, such as PowerInfer-2 (Xue et al., 2024), TurboSparse (Song et al., 2024c), ProSparse (Song et al., 2024a), and Q-Sparse (Wang et al., 2024b). These works require training on hundreds of billions of tokens and consume substantial hardware resources. There are also works, such as InfiniGen (Lee et al., 2024a), NSA (Yuan et al., 2025), and SeerAttention (Gao et al., 2024), focusing on KV cache sparsity but not weight sparsity. These methods benefit long context scenarios (¿32K) which are not the common cases on edge. TEAL (Liu et al., 2025a) proposes a training-free, magnitude-based sparsity method (see Fig. 3), where only activations above a threshold are computed. However, the active weights cannot be predicted, but only be identified after the input activation is ready. Additionally, the method is empirical, and there is no mechanism to compensate for the accuracy loss due to the potential misidentification of active weights. Therefore, current techniques fall short of achieving adaptive memory usage for LLMs.
@@
Firstly, Cross-layer active weight preloading. To address the sequential dependency issue of active weights with its input activation in order to enable computation and loading overlapping, we propose cross-layer active weight preloading. It creatively utilizes the current layer’s activation to pre-identify the next n layers’ active weights. It is based on the obeservation that due to the widely used residual connection, the activation magnitude distribution across layers share significant similarity (¿80% shown in Fig. 4(a)). For the active weights that missed by pre-loading, ActiveFlow loads on-demand when the actual activation is ready.
@@
[…]
@@ 4. Active Weight Swapping Pipeline
spsp sparsity of LLM
hrhr average hit rate of weight cache
sisi average similarity of cross-layer group
@@
NN Layer number of a cross-layer group
@@ 5. Self-Distillation for Top-K Sparse LLM
Even with superior quality compared to other sparsity techniques, Top-K activation sparsity still introduces an approximation in active weight selection, especially in high sparsity. Traditional methods such as supervised fine-tuning often fail to recover the performance of the model under high sparsity, as they cannot effectively capture nuanced weight distributions and activation patterns caused by sparsity, leading to a degradation of precision. To address this, we propose Top-K sparsity-aware self-distillation, an extension of quantization and fine-tuning pipelines. It preserves the efficiency benefits of sparsity while substantially reducing computational overhead and enhancing both accuracy and generalization. In practice, it improves performance with only a few to tens of GPU hours on a few thousand samples, and generalizes effectively across different sparsity levels.
Self distillation. As shown in Fig. 13, we maintain the outputs of the dense (teacher) and sparse (student) models, using the soft output distribution of the dense model as supervision. This allows the sparse student to capture richer correlations than hard labels and preserve fine-grained distributional details, which is crucial to compensate for information loss induced by Top-K activation sparsity.
KL loss. We adopt the Kullback-Leibler Divergence (KLD) loss to measure the discrepancy between the student and teacher distributions:
(10) $\mathcal{D}_{\text{KL}}(P_{T}\parallel P_{S})$ $=\sum_{i}P_{T}(i)\log\frac{P_{T}(i)}{P_{S}(i)}$
Minimizing $\mathcal{D}_{\text{KL}}$ encourages the sparse model to closely mimic the dense teacher’s distribution, preserving essential weight correlations and improving performance under high sparsity. Our framework is also orthogonal to quantization, as the KL-based distillation loss depends only on the output distribution and thus remains fully compatible with QAT, making it complementary to quantization errors and enabling additional efficiency gains with minimal accuracy loss.
Gradient STE. Gradient vanishing is a common issue when fine-tuning activation-sparse models, as the sparsity mask sets many elements to zero, preventing gradients from being properly propagated and slowing or even blocking convergence. To mitigate this, we employ gradient Straight-Through Estimation (STE), which replaces the gradient of the masking operation with an identity function during backpropagation:
@@
This allows gradients to flow as if the mask contains non-zero values, ensuring sufficient update signals even under high sparsity. Consequently, STE accelerates convergence, enhances training robustness, and helps the model preserve critical weight distributions and activation patterns.
Figure 13. The forward and backward of self-distillation between teacher and student.
Inherent adaptability. A key advantage of our self-distillation framework is its inherent adaptability across sparsity levels.
In conventional fine-tuning or distillation pipelines, models at different sparsity ratios typically require separate training processes, which is both time-consuming and computationally expensive.
In contrast, our method requires only a single distillation at a fix sparsity level, where the student is forced to maintain performance under extreme information constraints. The distilled model not only captures the dense teacher’s distributional characteristics but also develops robustness in reconstructing critical features, enabling direct inference across other sparsity levels without additional fine-tuning. This robustness naturally extends with increasing activation budget: model leverages the richer signals without retraining, achieving near-lossless performance over a wide sparsity range. We refer to this property as one-distill-all-scale; as Table 3 shows, PPL error remains within 1% even when training and inference sparsity differ by up to $\pm 15\%$.
By eliminating the need for repeated training at each sparsity level, this approach greatly reduces overall cost while ensuring consistent accuracy and efficiency in sparse LLM deployment.
@@ 6. Implementation
[…]
@@ 7. Evaluation
7.1. Evaluation setup
Measurement. Our evaluation focuses on decoding speed, perplexity, accuracy, latency, hit rate, memory cost, power, and energy consumption. We use the clock_gettime function to record start and end timestamps, computing latency as the difference between them. We measure the total number of decoded tokens and the total decoding time, calculating speed as $N_{tokens}/Latency$. We use lm-eval-harness, a widely used LLM evaluation framework, to measure perplexity on the WikiText-2 dataset and accuracy on five downstream tasks: 5-shot MMLU, 5-shot GSM8K, 25-shot ARC Challenge, 25-shot ARC Easy, and 0-shot PIQA. We track cache hits and misses, computing the hit rate as $N_{hit}/(N_{hit}+N_{miss})$. We analyze memory cost using the Android Studio Profiler. We obtaine current and voltage values by reading system files (voltage_now and current_now) to calculate power consumption. These values are collected every 0.5 seconds on average, and we use the decoding latency to compute the overall energy consumption.
@@ 7.2. End-to-end performance
Perplexity and Downstream tasks. Our method demonstrates that large language models can maintain low perplexity under significantly reduced memory costs, e.g., achieving performance comparable to the full-weight setting for LLaMA-2-7B and LLaMA-3-8B at only 60% memory usage in Fig. 14(b), and matching the Mixtral-8x7B baseline (24.6GB) with just 4.4GB. While perplexity increases under more aggressive sparsity, our self-distillation strategy effectively alleviates performance degradation, enabling consistent improvements over TEAL across five downstream tasks, with gains up to 10.98% at 70% sparsity and average improvements ranging from 2.64% to 10.21% across sparsity levels in Table 4.
ablation study for self-distillation. To validate the effectiveness of each component in our framework, we conducted ablation studies focusing on the gradient straight-through estimator (STE) and self-distillation techniques. We carried out experiments on the Llama-3-8B model, comparing performance under different configurations, including: 1. removing STE; 2. replacing self-distillation with full fine-tuning. As shown in Table 5, our framework’s components improve model performance across different sparsity levels.
different models for self-distillation. We evaluated the proposed self-distillation framework across diverse model architectures and visualized the results in equivalent memory–performance plots. The method demonstrates strong generality: from standard 7B models to highly compressed Qwen2.5-0.5B and sparse MoE architectures, it consistently delivers significant sparsity-driven performance gains. Unlike conventional compression, our approach ensures predictable accuracy loss while achieving strict acceleration. As illustrated in Figure 15, our results consistently lie on the Pareto frontier of equivalent memory and performance, underscoring the framework’s adaptability for practical deployment.
Table 3. End-to-End PPL results under Varying Sparsity Levels of Meta-LLaMA-3-8B (distillation under Fixed-Sparsity), using 4-bit quantization.
Method 0% 50% 60% 70% 80%
TOP-K 6.6836 8.1950 10.0121 15.9046 96.3015
Ours 7.4510 8.3216 10.2442 16.1081
Ours distill on 50% 7.4510 8.5625 11.8981 41.5303
Ours distill on 60% 7.4636 8.3216 10.8789 27.2720
Ours distill on 70% 7.8163 8.4440 10.2442 19.6981
Ours distill on 80% 9.3462 9.8854 11.1767 16.1081
Figure 15. Pareto frontier of actual runtime memory vs. PPL for TEAL and our self-distillation.
Table 4. PPL and Downstream Task Accuracy of LLaMA-3-8B.We use TEAL as our baseline TOP-K method.
Method PPL MMLU GSM8K ARC-C ARC-Easy PIQA
Origin (0%) 6.0874 65.16% 50.87% 54.95% 83.96% 80.74%
TOP-K (50%) 7.7762 59.21% 32.30% 49.32% 81.40% 78.45%
TOP-K (60%) 9.0042 51.67% 17.76% 45.56% 76.98% 76.01%
TOP-K (70%) 13.6816 36.53% 3.34% 33.70% 66.60% 70.30%
TOP-K (80%) 73.1400 25.69% 1.67% 21.08% 38.89% 56.80%
Ours (50%) 6.9677 61.41% 38.89% 52.13% 81.48% 79.98%
Ours (60%) 7.7935 57.01% 28.58% 49.57% 79.46% 77.80%
Ours (70%) 9.5079 47.51% 12.89% 41.64% 74.12% 77.31%
Ours (80%) 14.6401 29.40% 2.05% 32.34% 62.30% 69.10%
@@ 7.3. Technique breakdown
[…]
@@
Figure 18. The performance of task-level and context-level cache.
@@
Contextual caching policy. Fig. 18 compares context-level and task-level caches. On BoolQ, when token length=10, the context-level cache achieves a 77% hit rate, 13% higher than task-level. As length increases to 40, the hit rate slightly drops to 74% but still remains 10% higher. Across downstream tasks (Fig. 18b), task-level hit rate varies between 54–74%, while context-level consistently adapts, yielding an average 12% improvement.
Cache efficiency. As shown in Fig. 19, enlarging cache size significantly reduces flash access. With 50% cache, flash operations shrink to 18% of weights, giving a 5.2× reduction in memory access compared to full loading. Larger cache further improves hit rate but also increases memory footprint; therefore, we adjust cache size dynamically based on available device memory.
Figure 19. The rate of attention Q/K/V weights to load of LLaMA-2-7B model with 50% sparsity under various cache sizes.
@@
Table 5. Llama-3-8B ablation studies
Method PPL MMLU GSM8K ARC-C ARC-Easy PIQA
Ours (50%) 6.9677 61.41% 38.89% 52.13% 81.48% 79.98%
Ours-Distill (50%) 7.4872 59.67% 32.83% 50.94% 81.82% 78.89%
Ours-STE (50%) 7.0660 60.78% 37.45% 49.91% 81.65% 79.54%
Ours (60%) 7.7935 57.01% 28.58% 49.57% 79.46% 77.80%
Ours-Distill (60%) 8.2635 55.73% 24.64% 50.68% 80.39% 77.53%
Ours-STE (60%) 8.1517 55.27% 23.58% 47.78% 78.20% 77.91%
Ours (70%) 9.5079 47.51% 12.89% 41.64% 74.12% 77.31%
Ours-Distill (70%) 9.9149 45.29% 10.77% 42.41% 73.99% 75.41%
Ours-STE (70%) 11.1969 37.72% 3.87% 35.92% 68.69% 73.18%
Ours (80%) 14.6401 29.40% 2.05% 32.34% 62.30% 69.10%
Ours-Distill (80%) 37.7404 24.44% 1.90% 20.14% 36.95% 57.67%
Ours-STE (80%) 23.2097 25.96% 1.59% 22.87% 52.74% 64.85%